OpenStudy (anonymous):

sharon neds 64 credits. She only has 24. What percent does she have?

OpenStudy (anonymous):

24/64 = 0.375 = 37.5 %

OpenStudy (triciaal):

percent means per 100 doing a ratio (comparing to get the equivalent) 24/64 =x/100 where x is the unknown % solve for x